Its a algebraic fraction please help it with me.Question: s=1/2, t=4 and u =1/3.
1)5t+s+u

Answer by stanbon(75887)

Its an algebraic fraction please help me with it.
Question: s=1/2, t=4 and u =1/3.
1)5t+s+u
----
Ans: 5(4) + (1/2) + (1/3)
----
Least common denominator:: 6
--------
= 20 + 3/6 + 2/6
-----
= 20 5/6
--------------
Cheers,
Stan H.
-------------